### v1

@@ 6 Deeper knowledge
of the true prior variance matrix $\Sigma$, its largest $J\in\{0,1,\ldots,K\}$ eigenvalues, and the corresponding eigenvectors, but does not know the smallest (KJ)(K-J) eigenvalues or the corresponding eigenvectors. Intuitively, he knows how much the features $v_{1},\ldots,v_{K}$ influence the states $\theta_{1},\ldots,\theta_{K}$ overall, and he knows the $J$ most influential features and how influential they are, but he does not know the (KJ)(K-J) least influential features or how influential they are. He assumes the latter features are equally influential. That is, he assumes $\theta$ has prior variance
@@
is the mean of the smallest (KJ)(K-J) eigenvalues of $\Sigma$.242424 If $J=K$, then we define $\lambda_{K}^{(J)}\equiv\lambda_{K}$. The matrix (21) has the same trace as $\Sigma$ but (possibly) different eigenvalues; its $k^{\text{th}}$ largest eigenvalue
@@
independently of $J$. Likewise $\lambda_{K}^{(0)}=\overline{\lambda}$ by definition. Thus $\Sigma^{(0)}=\overline{\lambda}I_{K}$ is the prior variance matrix used by the “naïve” agent described in Section 3.2. The parameter $J$ interpolates between $\Sigma^{(0)}$ and $\Sigma^{(K)}=\Sigma$. It captures the “depth” of the agent’s conceptual knowledge: raising $J$ corresponds to knowing more of the structural features that influence states.252525 Since $\lambda_{1}\geq\cdots\geq\lambda_{K}$ (by assumption), there are non-increasing returns to knowing more features (i.e., increasing $J$), since each additional feature contributes a non-increasing share of the states’ prior variances. This captures the intuitive idea that the agent acquires conceptual knowledge by learning about the most influential features first. For example, he could gain knowledge by taking classes or reading textbooks that provide “high-level summaries” before “digging into the details.” Accordingly, we say the agent has “$J$-deep conceptual knowledge” if his prior on $\theta$ has variance $\Sigma^{(J)}$.
Suppose the agent has $J$-deep conceptual knowledge and collects an optimal sample. Then, by analogy to (15) and (16), this sample has rank
@@
where $\tau\equiv n\overline{\lambda}/\sigma_{u}^{2}$ is the precision index defined in Section 5. For example, letting $J=0$ yields the rank $R^{(0)}=K$ and value
$\pi^{(0)}=\frac{\overline{\lambda}\tau}{K+\tau}$
of an optimal sample collected by a naïve agent. We refer to the difference
@@
between $\pi^{(J)}$ and $\pi^{(0)}$ as “the value of $J$-deep conceptual knowledge.” We characterize the relationship between $R^{(J)}$} and $J$ in Lemma 4, and the relationship between $\Pi^{(J)}$ and $J$ in Theorem 3.
@@
for each $J\in\{0,\ldots,K\}$. This threshold is non-decreasing in $\tau$.
@@
    equals $\Pi$ when $J\geq{R^{*}}$.
Theorem 3 says that deeper knowledge is (weakly) more valuable. Intuitively, the more the agent knows about states’ features, the better he can design samples that provide information about the most influential features.
The value of $J$-deep conceptual knowledge is bounded above by the value $\Pi^{(K)}=\Pi$ of “full” knowledge. It attains this bound when $J\geq{R^{*}}$. Thus, the agent gains no additional value from deepening his knowledge beyond depth $R^{*}$}. This is because he ignores the least influential $(K-{R^{*}})$ features when collecting data (since $\delta_{k}^{*}=0$ for each $k>{R^{*}}$), so learning about those features does not change his optimal sample or the value of the information it contains.
@@
with $0<\alpha<1$. Then $\lambda_{1},\ldots,\lambda_{K}$ are strictly positive, have mean $\overline{\lambda}=1$, are constant in the limit as $\alpha\to 0$, and undergo a MPS as $\alpha$ rises.262626 For each $k\in\{1,\ldots,K\}$ we have $\lambda_{k}\to 1$ as $\alpha\to 0$ by L’Hôpital’s rule. Moreover, the partial sum $\sum_{j=1}^{k}\lambda_{k}$ $=\frac{K\mathopen{}\mathclose{{\left(1-(1-\alpha)^{k}}}\right)}{1-(1-\alpha)^{K}}$ is non-decreasing in $\alpha$ and is constant in $\alpha$ when $k=K$. Thus, by Lemma 3, the eigenvalues $\lambda_{1},\ldots,\lambda_{K}$ undergo a MPS when $\alpha$ rises. This parameter determines the rate
@@
at which $\lambda_{k}$ decays as $k$ grows. Intuitively, the larger is $\alpha$, the faster features’ marginal influences on states diminishes. Thus, if $\alpha$ is larger, then states are more reducible.
@@
Figure 2 shows how $R^{(J)}$} and $\Pi^{(J)}$ depend on $J$ when $(K,\overline{\lambda},\sigma_{u}^{2})=(100,1,1)$ and $\lambda_{k}$ decays at rate $\alpha\in\{0.01,0.02,0.03\}$. If $J$ is sufficiently small, then the agent collects an optimal sample with full rank $R^{(J)}=100$; otherwise, he collects a sample with rank ${R^{(J)}}=\min\{J,{R^{*}}\}$. The threshold depth at which he switches from 100 to $\min\{J,{R^{*}}\}$ rises as the sample size $n$ rises, consistent with Lemma 4.272727 Here $\tau=n$ because $\overline{\lambda}=\sigma_{u}^{2}=1$. Intuitively, if the agent has access to more data, then he has less to gain from knowing which features are most influential and focusing on them when he collects data.
The value $\Pi^{(J)}$ is increasing in $J$ when $J<{R^{*}}$ and constant when $J\geq{R^{*}}$, consistent with Theorem 3. It is increasing in $\alpha$, consistent with Theorem 1: conceptual knowledge is more valuable when the eigenvalues $\lambda_{1},\ldots,\lambda_{K}$ are more spread out. Likewise $\Pi^{(J)}$ is non-monotone in $n$, consistent with Theorem 2: raising $n$ allows the agent to learn more about the “in-sample” features (raising $\Pi^{(J)}$) but also prompts him to expand his sample to include more features (lowering $\Pi^{(J)}$).
7 Deeper knowledge or more data?
Finally, we consider the trade-off between deepening the agent’s knowledge and giving him access to more data.282828 In contrast, Dominitz and Manski, (2017) study the trade-off between having more data and “better” data. Suppose he has $J$-deep knowledge and collects an optimal sample of size $n$. The value $\pi^{(J)}$ of this sample indexes his welfare: it is larger when his minimized posterior expected loss is smaller. Lemma 5 says he is better off with deeper knowledge and more data. Intuitively, if he has deeper knowledge, then he can use it to design “better” samples that provide information about the “right” features. If he has access to more data, then he can obtain more information, making his posterior beliefs more accurate and his expected loss lower.
@@ 7 Deeper knowledge or more data?
The value $\pi^{(J)}$ is
@@
be the minimum sample size necessary to attain this value. This size is smaller when the agent has deeper knowledge and when states are more reducible:
@@
Fix $\pi_{0}\geq 0$. Then $n_{\pi_{0}}^{(J)}$
@@
    is non-increasing in $J$ and
@@
    does not rise when $\lambda_{1},\ldots,\lambda_{K}$ undergo a MPS.
Theorem 4 says that if the agent has deeper knowledge or states are more reducible (holding the depth of his knowledge constant), then he can attain the same welfare with less data. This is because he can design better samples and extract more value from each observation, lowering the number he needs to attain the target $\pi_{0}$.
@@
As in illustration of Theorem 4, suppose the “true” eigenvalues $\lambda_{1},\ldots,\lambda_{K}$ have mean $\overline{\lambda}=1$ and decay at rate $\alpha$ as in Section 6. Figure 3 shows how the minimum sample size $n_{\pi_{0}}^{(J)}$ depends on the welfare target $\pi_{0}$ and depth $J$ when $(K,\sigma_{u}^{2})=(100,1)$ and $\alpha\in\{0.01,0.02,0.03\}$. Given $\pi_{0}$, the size $n_{\pi_{0}}^{(J)}$ is decreasing in $J$ when $J<{R^{*}}$ and constant in $J$ when $J\geq{R^{*}}$. Intuitively, if the agent’s knowledge is too shallow, then it constrains his ability to design samples that provide information about the “right” features. Deepening his knowledge relaxes this constraint. It empowers him to design better samples, extract more value from each observation, and require fewer observations to attain $\pi_{0}$. However, once his knowledge is deep enough, deepening it further does not change how he designs samples or the marginal value of each observation. So the only way to collect more valuable data is to collect more data, thus making $n_{\pi_{0}}^{(J)}$ constant in $J\geq{R^{*}}$.
The curves in Figure 3 are indifference curves: they trace out sets of depth-size pairs $(J,n)$ that allow the agent to attain different welfare targets $\pi_{0}$. The slope of each curve equals the marginal rate of substitution (hereafter “MRS”) between knowledge and data. Intuitively, this MRS captures the number of observations an additional unit of conceptual knowledge is “worth.” It depends on the depth-size pair $(J,n)$, the target $\pi_{0}$, and the parameter $\alpha$ indexing states’ reducibility. Raising this parameter raises the rank $R^{*}$} of the optimal sample he would collect if he had full knowledge (see Figure 2). So raising $\alpha$ can have three effects on the MRS between knowledge and data:
@@
Thus, if states are more reducible, then the MRS between knowledge and data may can be higher or lower, depending on the depth of the agent’s knowledge.
Overall, Figure 3 suggests that if the agent has deeper knowledge, then he can attain the same welfare with less data, especially when states are highly reducible. However, this pattern may be specific to the case in which $\lambda_{1},\ldots,\lambda_{K}$ decay at the constant rate $\alpha$. We defer analyzing the general case (in which $\lambda_{1},\ldots,\lambda_{K}$ are distributed arbitrarily) to future research.
@@ 8 Related literature
We assume states and outcomes are jointly normally distributed under the agent’s prior, and his actions are real-valued and induce quadratic losses. This assumption allows us to derive closed-form expressions for the values of information and conceptual knowledge. It is common in the literature on statistical decisions (Hastie et al.,, 2009), and in the literature on learning and information acquisition. Within the latter literature, two closely related papers are by Whitmeyer, (2024) and Ilut and Valchev, (2025).
Whitmeyer, (2024) studies the value of information in abstract decision problems. His Theorem 3.1 says that information is more valuable when an agent’s optimal payoff is more convex in their beliefs. This result is analogous to our Theorem 1: an increase in convexity is analogous to eigenvalues undergoing a MPS.
Ilut and Valchev, (2025) consider an agent who learns a function via “abstract reasoning” and “integrating experience.” Similarly, we consider an agent who learns a vector by using conceptual knowledge and observing data. Ilut and Valchev, study a dynamic setting, and focus on the “learning traps” that arise from reasoning too little or having the wrong data. In contrast, we study a static setting, and focus on the benefits of reasoning correctly and having the “right” data.
Ilut and Valchev, model the cognitive processes that humans use to learn and make decisions. These processes differ from those used by machines: whereas humans can use concepts and causal reasoning, machines currently cannot (without human supervision). Instead they rely on pattern recognition and data-driven prediction (Felin and Holweg,, 2024).
This difference between humans and machines motivates papers comparing their predictive performance (e.g., Kleinberg et al.,, 2018; Kühl et al.,, 2022; Mullainathan and Obermeyer,, 2022). We shed light on when humans are likely to outperform “naïve” (in the sense defined in Section 3.2) machines: when the unknowns are highly reducible (see Theorem 1), when data are noisy or scarce (see Theorem 2), and when the sampling frame is limited (see Appendix Section A2.3).
We also identify one reason why humans may outperform machines: humans have structural knowledge that empowers us to “ask the right questions.” This knowledge tells us which features of an environment are most important, so we can focus on them when we collect and analyze data. We do not claim this idea as our own: Fessler and Kasy, (2019) discuss how structural knowledge can be used to improve econometric estimators; Jackson, (2019) discusses how it can guide our experimental designs. However, as far as we know, we are the first to quantify the value of having structural knowledge when learning and making decisions.
Whereas we compare humans and machines implicitly, Iakovlev and Liang, (2025) compare them explicitly. They study the “value of context”: how much predictive power one gains from choosing the “right” covariates, relative to an algorithm that cannot make this choice. They show that the value of context vanishes as one’s access to data becomes large. Similarly, we show that the value of conceptual knowledge vanishes as one’s access to data becomes large (see Theorem 2). Together, these results suggest that knowledge and data are substitutes in “big data worlds”: machines can use large, rich samples to learn and make decisions without prior structural knowledge. However, our Theorem 2 also suggests that knowledge and data are complements in “small data worlds”: humans can make small samples more valuable by leveraging our structural knowledge.
This paper also connects to the literature on model-based learning. Andrews et al., (2025) find that “black box” algorithms outperform models when predicting within domains, but are worse at generalizing across domains.292929 See also Fudenberg and Liang, (2019), Peterson et al., (2021), and Peysakhovich and Naecker, (2017) for comparisons of model-based and black box predictions. This is consistent with our idea that models embed conceptual knowledge (see Appendix Section A1.2) and that such knowledge boosts out-of-sample predictive performance (see Appendix Section A2.3). Fudenberg et al., (2022) propose a measure of model “completeness”: the share of reducible prediction error that imposing a model reduces. Our paper suggests a different notion of completeness: a model is “more complete” when it embeds deeper conceptual knowledge.303030 Mailath and Samuelson, (2020) argue that “in practice, people work with models that are deliberately incomplete, including the most salient variables and excluding others.” Indeed, Theorem 3 implies that the agent does not benefit from using models embedding depths greater than the rank $R^{*}$} of an optimal sample.
In our analysis of deeper knowledge (see Sections 6 and 7), we assume there exists a “true” model of the agent’s environment (i.e., a true structural relationship among the unknown states) that can be known at different depths. This is in contrast to the literature on model uncertainty and mis-specification, which considers settings where the true model is unknown (e.g., Cerreia-Vioglio et al.,, 2025; Esponda and Pouzo,, 2016; Hansen and Sargent,, 2001; Marinacci,, 2015).313131 See also Chatfield, (1995) for a discussion of model uncertainty in statistics. In these settings, many authors advocate a “robust” approach that imposes minimal structural restrictions and provides payoff guarantees across a range of possible models (Gilboa and Schmeidler,, 1989; Klibanoff et al.,, 2005). This is consistent with our “naïve” agent, who lacks conceptual knowledge, assuming a minimally restrictive prior. We add to the model uncertainty and mis-specification literature by quantifying the value of knowing and imposing the correct structural restrictions.
@@
We also assume away any competitive or political forces that may lead the agent to choose one model over another (Dasaratha et al.,, 2025; Izzo et al.,, 2023). These forces arise in the literature on “models as narratives” (Aina,, 2025; Schwartzstein and Sunderam,, 2021; Eliaz et al.,, 2025). Papers in that literature focus on settings where many models are plausibly “true” (e.g., financial markets and political campaigns). In contrast, many real-world decisions are made in settings where there is an objectively true model one could uncover with enough education and introspection.323232 For example, there exists an objectively true set of mechanisms through which fertilizers help plants grow. A farmer could discover these mechanisms experimentally or be taught them (see Sankar et al.,, 2025), or may rely on heuristics that capture some mechanisms but not all (e.g., the farmer may notice that two fertilizers intensify greening, but not know they do so because they supply nitrogen). As another example, a physics student may learn the objectively true model $E=mc^{2}$, then deepen his knowledge by learning about Lorentz factors. Such settings are our focus in this paper. For this reason, we do not consider the issues that would arise if the agent’s conceptual knowledge was mis-specified or supplied by a strategic communicator with competing incentives. However, we believe these issues are interesting and worthy of future research.

### v2

@@ 6 Deeper conceptual knowledge
of the true prior variance matrix $\Sigma$ and its first $J\in\{0,1,\ldots,K\}$ eigenpairs $(\lambda_{k},v_{k})$, but does not know the last (KJ)(K-J) eigenpairs. Intuitively, he knows the $J$ concepts with the most explanatory power, but does not know the (KJ)(K-J) concepts with the least explanatory power. So he assumes the components of $\theta$ orthogonal to $v_{1},\ldots,v_{J}$ have equal prior variances; specifically, he assumes $\theta$ has prior variance
@@
is the mean of the smallest (KJ)(K-J) eigenvalues of $\Sigma$. [^32] If $J=K$, then the bracketed term in (26) equals zero and so $\lambda_{K}^{(K)}$ can be defined arbitrarily. We define $\lambda_{K}^{(K)}\equiv\lambda_{K}$. The matrix (26) has the same trace as $\Sigma$ but (possibly) different eigenvalues; its $k^{\text{th}}$ largest eigenvalue
@@
independently of $J$. Likewise $\lambda_{K}^{(0)}=\overline{\lambda}$ by definition. Thus $\Sigma^{(0)}=\overline{\lambda}I_{K}$ is the naïve prior variance matrix discussed in Sections 3.2 and 5. The parameter $J$ interpolates between $\Sigma^{(0)}$ and $\Sigma^{(K)}=\Sigma$. It captures the “depth” of the agent’s conceptual knowledge: the larger is $J$, the more concepts he knows and the richer is his mental model of $\theta$. [^33] Since $\lambda_{1}\geq\cdots\geq\lambda_{K}$ (by assumption), there are non-increasing returns to knowing more concepts (i.e., increasing $J$), since each additional concept contributes a non-increasing share of states’ prior variances. Intuitively, when the agent acquires conceptual knowledge, he prioritizes concepts with more explanatory power. For example, he takes classes or reads textbooks that provide “high-level summaries” before “digging into the details.” , [^34] For example, a farmer could deepen his conceptual knowledge by learning about different nutrients that crops need and fertilizers supply (e.g., nitrogen, phosphorus, and potassium). He could also learn about the nuances of nutrients he already knows about. Nutrients can take different chemical forms (e.g., nitrogen can take the form of ammonium and nitrate) that have different effects when applied to different crops. Learning about these differences gives the farmer new concepts that explain fertilizers’ overall effects.
We say the agent has “$J$-deep conceptual knowledge” if his prior on $\theta$ has variance $\Sigma^{(J)}$. Suppose he has such knowledge and designs an optimal sample. Then, by analogy to (19) and (20), this sample has rank
@@
where $\tau\equiv n\overline{\lambda}/\sigma_{u}^{2}$ is the precision index defined in Section 5. For example, letting $J=0$ yields the rank $R^{(0)}$ and value $\pi^{(0)}$ of an optimal sample collected by a naïve agent. We refer to the difference
@@
between $\pi^{(J)}$ and $\pi^{(0)}$ as the “value of $J$-deep conceptual knowledge.” We characterize the relationship between $R^{(J)}$} and $J$ in Lemma 4, and the relationship between $\Pi^{(J)}$ and $J$ in Theorem 3.
@@
for each $J\in\{0,\ldots,K\}$. This threshold is non-decreasing in the precision index $\tau$.
@@
    equals the value $\Pi$ of full knowledge when $J\geq{R^{*}}$.
Theorem 3 says that deeper knowledge is (weakly) more valuable. Intuitively, knowing more concepts allows the agent to design samples that provide more payoff-relevant information.
The value of $J$-deep conceptual knowledge is bounded above by the value $\Pi^{(K)}=\Pi$ of “full” knowledge, and attains this bound when $J\geq{R^{*}}$. Thus, the agent gains no additional value from knowing more than the $R^{*}$} concepts with the most explanatory power. This is because he ignores the other $(K-{R^{*}})$ concepts when he designs samples (since $\delta_{k}^{*}=0$ for each $k>{R^{*}}$), so knowing those concepts does not change his optimal sample.
@@
with $0<\alpha<1$. Then $\lambda_{1},\ldots,\lambda_{K}$ are strictly positive, have mean $\overline{\lambda}=1$, are constant in the limit as $\alpha\to 0$, and undergo a MPS as $\alpha$ rises. [^35] For each $k\in\{1,\ldots,K\}$ we have $\lambda_{k}\to 1$ as $\alpha\to 0$ by L’Hôpital’s rule. Moreover, the partial sum $\sum_{j=1}^{k}\lambda_{k}$ $=\frac{K\mathopen{}\mathclose{{\left(1-(1-\alpha)^{k}}}\right)}{1-(1-\alpha)^{K}}$ is non-decreasing in $\alpha$ and is constant in $\alpha$ when $k=K$. Thus, by Lemma 3, the eigenvalues $\lambda_{1},\ldots,\lambda_{K}$ undergo a MPS when $\alpha$ rises. This parameter determines the rate
@@
at which $\lambda_{k}$ decays as $k$ grows. Intuitively, the larger is $\alpha$, the faster concepts’ marginal explanatory power falls. Thus, if $\alpha$ is larger, then states are more reducible.
@@
Figure 3 shows how $R^{(J)}$} and $\Pi^{(J)}$ depend on $J$ when $(K,\overline{\lambda},\sigma_{u}^{2})=(100,1,1)$ and $\lambda_{k}$ decays at rate $\alpha\in\{0.01,0.02,0.03\}$. If $J$ is sufficiently small, then the agent designs an optimal sample with rank $R^{(J)}=100$; otherwise, he designs a sample with rank ${R^{(J)}}=\min\{J,{R^{*}}\}$. The threshold depth at which he switches from 100 to $\min\{J,{R^{*}}\}$ rises as the sample size $n$ rises, consistent with Lemma 4.363636Here $\tau=n$ because $\overline{\lambda}=\sigma_{u}^{2}=1$. The value $\Pi^{(J)}$ of $J$-deep conceptual knowledge is increasing in $J$ when $J<{R^{*}}$ and constant in $J$ when $J\geq{R^{*}}$. This value is increasing in $\alpha$, consistent with Theorem 1: conceptual knowledge is more valuable when $\lambda_{1},\ldots,\lambda_{K}$ are more spread out. Likewise $\Pi^{(J)}$ is non-monotone in $n$, consistent with Theorem 2: raising $n$ allows the agent to learn more about the “in-sample” coefficients $\gamma_{1},\ldots,\gamma_{R^{(J)}}$ (raising $\Pi^{(J)}$), but also prompts him to expand his sample and learn about more coefficients (lowering $\Pi^{(J)}$).
7 More concepts or more data?
Finally, we study the trade-off between conceptual and statistical knowledge. Our model in Section 6 provides a formal language for describing this trade-off: would the agent rather know more concepts (i.e., increase the depth $J$) or have more data (i.e., increase the sample size $n$)?
Suppose the agent has $J$-deep conceptual knowledge and designs an optimal sample of size $n$. The value $\pi^{(J)}$ of this sample indexes the agent’s welfare: it is larger when his minimized posterior expected loss is smaller. Lemma 5 says the agent is better off with deeper knowledge or more data. Intuitively, if he knows more concepts, then he can “ask better questions.” If he has more data, then he can obtain better answers to his questions, making his posterior beliefs more precise and his expected loss lower.
@@ 7 More concepts or more data?
The value $\pi^{(J)}$ of an optimal sample designed by an agent with $J$-deep conceptual knowledge is
@@
be the minimum sample size necessary to attain this value. This size is smaller when the agent knows more concepts and when states are more reducible:
@@
Fix $\pi_{0}\geq 0$. The minimum sample size $n_{\pi_{0}}^{(J)}$ necessary to design a sample with value $\pi_{0}$
@@
    is non-increasing in the depth $J$ of the agent’s conceptual knowledge and
@@
    does not rise when the eigenvalues $\lambda_{1},\ldots,\lambda_{K}$ undergo a mean-preserving spread.
Theorem 4 says that if the agent knows more concepts, then he can attain the same welfare with fewer observations, especially when states are highly reducible. This is because he can design better samples and extract more value from each observation, lowering the number he needs to attain the target $\pi_{0}$.
@@
As in illustration of Theorem 4, suppose the true eigenvalues $\lambda_{1},\ldots,\lambda_{K}$ have mean $\overline{\lambda}=1$ and decay at rate $\alpha$ as in Section 6. Figure 4 shows how the minimum sample size $n_{\pi_{0}}^{(J)}$ depends on the welfare target $\pi_{0}$ and depth $J$ when $(K,\sigma_{u}^{2})=(100,1)$ and $\alpha\in\{0.01,0.02,0.03\}$. Given $\pi_{0}$, the size $n_{\pi_{0}}^{(J)}$ is decreasing in $J$ when $J<{R^{*}}$ and constant in $J$ when $J\geq{R^{*}}$. Intuitively, if the agent knows too few concepts, then he cannot design samples that focus on all of the “right” concepts. Giving him more concepts empowers him to design better samples, extract more value from each observation, and require fewer observations to attain $\pi_{0}$. However, once he knows all the “right” concepts, giving him more does not change how he designs samples or the marginal value of each observation. Then the only way to for him to obtain more valuable data is to obtain more data, thus making $n_{\pi_{0}}^{(J)}$ constant in $J\geq{R^{*}}$.
The curves in Figure 4 are indifference curves: they trace out sets of depth-size pairs $(J,n)$ that allow the agent to attain different welfare targets $\pi_{0}$. The slope of each curve equals the marginal rate of substitution (hereafter “MRS”) between concepts and data. Intuitively, this MRS captures the number of observations the agent would give up to know another concept. It depends on the depth-size pair $(J,n)$, the target $\pi_{0}$, and the parameter $\alpha$ indexing states’ reducibility. Raising this parameter raises the rank $R^{*}$} of the optimal sample he would design if he had full knowledge (see Figure 3). So raising $\alpha$ can have three effects on the MRS between concepts and data:
@@
So if states are more reducible, then the MRS between concepts and data may be higher or lower, depending on how many concepts the agent already knows. We defer analyzing this dependence to future research.
@@ 8 Related literature
Value of information
A large literature studies the instrumental value of information. Seminal work by Blackwell, (1951, 1953) shows that information sources can be ordered: one is more informative than another if it produces a sufficient statistic for the other’s information. Howard, (1966) and Raiffa and Schlaifer, (1961) link informativeness to instrumental value, defining the “value of information” as the difference in expected payoffs with and without it. More recently, Brooks et al., (2024) extend Blackwell’s order to settings with multiple sources, Frankel and Kamenica, (2019) characterize measures of information’s value, and Whitmeyer, (2025) studies value-increasing transformations in abstract decision problems. [^37] Others study information’s value in specific decision problems. For example, Lehmann, (1988) and Athey and Levin, (2018) study its value in monotone decision problems, Persico, (2000) studies its value in auctions, and Cabrales et al., (2013) study its value in investment decisions.
Our contribution is to illustrate why some information is more valuable than others, focusing on a specific decision problem relevant to economists and statisticians. We show that the value of a sample depends on its alignment with the agent’s conceptual structure: the prior variance matrix. Proposition 1 provides sharp bounds: a sample is most (least) valuable when the eigenvectors of the induced Gram and prior variance matrices are maximally aligned (mis-aligned). Proposition 2 characterizes the design of maximally valuable samples, showing that they focus on dimensions with the most prior variance. Theorem 1 quantifies how much this focus raises samples’ value.
We separate samples’ design (i.e., the covariates they contain) from their quality (i.e., signals’ precision). Blackwell, and his successors take samples as given, and ask how valuable they are. In contrast, we show how one can use prior knowledge to make samples more valuable. Thus, our framework allows us to transform the abstract notion of “informativeness” into concrete experimental design guidance.
Model-based inference
Another literature emphasizes the importance of models for interpreting data. This literature includes arguments by statisticians (e.g., Cox,, 1990) and computer scientists (e.g., Wolpert,, 1996) that models are essential for inference. In economics, Lucas, (1976) argues that models are necessary to make counterfactual predictions; Koopmans, (1947) and Wolpin, (2013) make similar arguments. Manski, (2003) shows formally that data alone are insufficient for inference: one must also make assumptions about the data-generating process (i.e., impose a model).
Recognizing that models are important, many authors study the consequences of not knowing the true model or using the wrong model. Gilboa and Schmeidler, (1989), Klibanoff et al., (2005), and Marinacci, (2015) develop decision theories under model uncertainty, while Cerreia-Vioglio et al., (2025) and Spiegler, (2016) develop decision theories under potential mis-specification.
@@
Others study the competitive and political forces that shape the models agents use (Dasaratha et al.,, 2025; Izzo et al.,, 2023; Montiel Olea and Prat,, 2025). These forces arise in the literature that treats models as “narratives”: stories that explain observed events (Aina,, 2025; Eliaz et al.,, 2025; Levy et al.,, 2022; Schwartzstein and Sunderam,, 2021). [^38] Papers in the “narratives” literature focus on settings where many models are plausibly correct (e.g., financial markets and political campaigns). In contrast, many real-world decisions are made in settings where there is a single, objectively correct model that can be learned through education and introspection. For example, there is an objectively correct set of mechanisms through which fertilizers help crops grow, which a farmer could discover experimentally or be taught as in Sankar et al., (2025). Such settings are our focus in this paper.
In contrast, we treat models as dimension-reduction devices. In our framework, a model specifies (i) which components of the states vary together (i.e., the eigenvectors $v_{1},\ldots,v_{K}$) and (ii) how much variance each component contributes (i.e., the eigenvalues $\lambda_{1},\ldots,\lambda_{K}$). The distribution of these variances determines how much the agent’s model reduces the $K$-dimensional state space to a lower-dimensional subspace. Theorem 1 characterizes the instrumental value of this reduction.
We also formalize what it means to have a “richer” model: knowing more eigenpairs of the prior variance matrix (see Section 6). We introduce a parameter $J$ that interpolates between having a diffuse model ($J=0$), a partial model ($J<K$), and a complete model (J=KJ=K). [^39] Thus $J$ offers an alternative to the “completeness” measure proposed by Fudenberg et al., (2022). , [^40] Mailath and Samuelson, (2020) argue that “in practice, people work with models that are deliberately incomplete, including the most salient variables and excluding others.” Indeed, Theorem 3 implies that the agent does not benefit from using models embedding depths greater than the rank $R^{*}$} of an optimal sample. Introducing this parameter allows us to study the trade-off between having a richer model (i.e., increasing $J$) or more data (i.e., increasing the sample size nn). [^41] In contrast, Dominitz and Manski, (2017) study the trade-off between having more data or “better” data. Theorem 4 characterizes this trade-off.
Thus, our analysis contributes by describing why models are important and quantifying how important they are. Our approach formalizes Jackson,’s (2019) claim that models are useful because they guide experimental design: our agent uses his model to design more valuable samples.
Cognition
A third literature studies the cognitive foundations of human learning. Johnson-Laird, (1983) argues that mental models are core inferential tools. Murphy, (2002) explains how humans build models from concepts. Tenenbaum et al., (2011) and Mitchell, (2021) discuss how concepts help humans generalize. Felin and Holweg, (2024) discuss how humans generate novel insights via directed experimentation.
We present a formal economic model that embeds these ideas. Our agent has a mental model of states as combinations of concepts. These concepts allow him to generalize: he can use signals of one state to make inferences about another. They also facilitate directed experimentation: knowing more concepts allows the agent to design more informative samples (see Theorems 3 and 4).
Ilut and Valchev, (2025) present a related but distinct model. They isolate two learning modes—“abstract reasoning” and “integrating experience”—that correspond to our notions of conceptual and statistical knowledge. Ilut and Valchev, study a dynamic setting, and focus on the “learning traps” that arise from reasoning too little or having the wrong data. In contrast, we study a static setting, and focus on the benefits of reasoning correctly and having the “right” data.
For our agent, the “right” data focus on an endogenous number of high-variance dimensions. This form of selective attention is optimal given his decision problem and learning environment. In contrast, Gabaix, (2014) and Schwartzstein, (2014) take selective attention as given, and study its consequences for decision-making and belief updating. Kőszegi and Szeidl, (2013), like us, study an agent who focuses on high-variance dimensions. But Kőszegi and Szeidl, assume such behavior as part of their model, whereas we derive it as a consequence of our model.
Human-machine comparisons
Finally, our paper connects to the literature comparing humans and machines. Tenenbaum et al., (2011) discuss how humans use concepts to learn from limited data, whereas machines rely on recognizing patterns in large sets of data (Goodfellow et al.,, 2016; Halevy et al.,, 2009). Our framework formalizes this distinction: conceptual knowledge compensates for data scarcity (see Theorem 4) but loses its value when data are abundant (see Theorem 2). Moreover, we identify a mechanism through which humans learn faster than machines: humans have conceptual knowledge that empowers them to “ask the right questions.”
Our findings parallel Iakovlev and Liang,’s (2025). They study the “value of context”: the gain in predictive performance from choosing the “right” model, rather than relying on pattern recognition. Iakovlev and Liang, show that the value of context vanishes with abundant data. Likewise, the value of conceptual knowledge vanishes with abundant data. However, our framework offers additional insight: the value of conceptual knowledge rises before it falls, changing direction at the threshold $\tau^{\prime}$ identified in Theorem 2.
Empirical work compares human and machine predictions in many domains, such as crime (Berk,, 2017; Kleinberg et al.,, 2018), finance (Jansen et al.,, 2025), and healthcare (Agarwal et al.,, 2023; Mullainathan and Obermeyer,, 2022). A common finding is that human predictions are less accurate because they are based on intuitive, potentially biased judgments, whereas machine predictions are based on statistical models (Angelova et al.,, 2025; Hoffman et al.,, 2018). Our agent’s predictions are also based on a model: his prior on the state vector. We study the gain from using a restrictive model (i.e., imposing the true prior $\mathbb{P}$) rather than a diffuse model (i.e., imposing the naïve prior $\mathbb{P}^{(0)}$).
Finally, Andrews et al., (2025) study model-based predictions with those made by “black box” algorithms, finding that models dominate at out-of-domain transfer. This empirical result supports one of our theoretical claims: models embed conceptual knowledge that improves out-of-sample predictive performance (see Appendix Section A2.3).